The basic WACC equation

The calculation of WACC involves calculating the weighted average of the required rates of return on debt, preferred stock, and common equity, where the weights equal the percentage of each type of financing in the firm’s overall capital structure.

is the symbol that represents the before-tax cost of debt in the weighted average cost of capital (WACC) equation.

Wyle Co. has $2.3 million of debt, $2 million of preferred stock, and $2.1 million of common equity. What would be its weight on common equity?
Answer #1

Total value of capital structure = 2.3 +2 + 2.1 = 6.4

Weight of common equity = 2.1 / 6.4

Weight of common equity = 0.33